OpenStudy (anonymous):

How much heat is lost by 2.0 grams of water if the temperature drops from 31oC to 29oC? The specific heat of water is 4.184 J/g•oC.

OpenStudy (anonymous):

A. 4.0 J B. 6.2 J C. 8.4 J D. 16.7 J

OpenStudy (shamim):

i thinku know the equation \[H=ms(\theta _{2}- \theta _{1})\]

OpenStudy (shamim):

here H= heat loss by water m= mass of water s= specific heat of water

OpenStudy (anonymous):

appreciated shamim

OpenStudy (shamim):

and \[\theta _{1}= \lower temperature= 29\]

OpenStudy (shamim):

so u got it

OpenStudy (shamim):

is not it

OpenStudy (dls):

\[\LARGE Heat~ Lost=-mc(T_2-T_1)\] \[\LARGE Heat Lost=mc(T_1-T_2)\](I removed the negative by reversing the brackets) Substitute the values.

OpenStudy (anonymous):

thanks shaminm n dls, this would help

OpenStudy (shamim):

welcome

OpenStudy (dls):

:)